Dissolve 500mg calcium carbonate
tablet in 20 mL distilled water and 20.00 mL 2 M HCl. Transfer each solution quantitatively with distilled water to a 250 mL volumetric flask. Dilute to volume and mix.
Titrate 15.00mL aliquots of the tablet solutions with 0.1 NaOH to the phenolphthalein endpoint.
Please determine the expected titration volume of tablet. thank you!
Answer #1

The first reaction that occurs is:

CaCO3 + 2 HCl = CaCl2 + CO2 + H2O

The moles of CaCO3 and HCl used are calculated:

n CaCO3 = g / MM = 0.5 g / 100 g / mol = 0.005 mol

n HCl = M * V = 2 M * 0.02 L = 0.04 mol

The moles of HCl that should remain in 250 mL solution are calculated:

n HCl rem = 0.04 - (0.005 * 2) = 0.03 mol

The moles of HCl in the 15 mL aliquot are calculated:

n HCl = 0.03 * 15/250 = 0.0018 mol

The volume of NaOH required to neutralize the excess HCl is calculated:

V NaOH = n * 1000 / M = 0.0018 * 1000 / 0.1 = 18 mL

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.